Fermi National Laboratory


Accelerator Update

Wednesday May 5
The day shift began with Operations monitoring store 3469, the Recycler stash, and the Antiproton Source (Pbar) stack, with the Main Injector (MI) down due to a power supply problem, and with MiniBooNE taking beam.

At 9:31 AM, a Pbar expert began a reverse proton study.

Operations reported at 11:19 AM that a Booster RF station (BRF16) had a short on its high voltage system. Operators bypassed the station.

Pbar resumed stacking at 11:56 AM.

A Recycler expert began an antiproton to MI study at 1:12 PM. This study halted stacking.

Stacking resumed at 2:12 PM.

A MI expert conducted a 1.5-hour barrier bucket study.

At 8:13 PM, Recycler experts calibrated their emittance monitors. This calibration halted stacking.

The MI RF area suffered a vacuum burst at 10:08 PM. Everything reset okay.

Thursday May 6
The midnight shift began with Operations monitoring store 3469, with Pbar stacking, and with MiniBooNE taking beam.

At 6 AM, Operations halted all beam, except for the store, in preparation for maintenance work accesses in Booster, MI, and Linac.

Tevatron (TeV) experts began an end-of-store study at 12:27 PM.

Linac came out of access and Operations reestablished beam to the Linac dump at 12:43 PM.

Booster came out of access and Operations reestablished beam to the Booster dump at 13:40 PM.

Operations reported, at 3:52 PM, a problem with a MI Lambertson (I:LAM10). Techs had the Lambertson repaired by 6:40 PM and beam to the Accumulator by 6:47 PM

Operations established beam to MiniBooNE at 7:20 PM.

Operations terminated store 3469 at 8:22 PM.

Friday May 7
The midnight shift began with the TeV in shot setup, with operations monitoring the stash, and with MiniBooNE taking beam.

At 12:26 AM, Operations established store 3471 with and initial luminosity of 54.7E30. Pbar resumed stacking.

Operations reported losing the Booster Interlock permit at 5:45 AM due to a bad vacuum valve status. Operators called in an expert. This halted all down stream beam.

The Booster permit came back and Operations resumed sending beam to Pbar at 6:54 AM.

Plans for Today and Weekend
The plans for today and this weekend are to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/7/2004   email Fermilab